OpenStudy (anonymous):

x^3y = 27 what does x^y equal? I solved for y and got y = 9 but I am unsure how to get x. Possible answers are: 3,6,9,12,81

jimthompson5910 (jim_thompson5910):

Hint: x^(3y) = 27 (x^y)^3 = 27 (x^y)^3 = (3)^3